OpenStudy (anonymous):

The number of male students at the local college increased by 250 more male students than the year before. Given that the number of males students that attended college the year before was 25,000, determine the growth factor. a. 1.01 b. 1.10 c. 1.001 d. 1.11

OpenStudy (anonymous):

@Elise_a18 @RECKLESS43 @TuringTest @aryandecoolest

OpenStudy (anonymous):

umm ok so U knw Growth Factor can be defined as the ratio of the change in the number of things/people, divided by the original number of things/people.

OpenStudy (anonymous):

Right so, In this case the change in number equals: 250. Original Number equals: 25,000. Growth Factor = 250/25000 = 0.001( Not in option) @Rosepetal_